Tuesday 09/11/12

9:10 AM – 7:15 P, 9:00P – 11:55 P

  • Log off Tuck (Reserved for Yari)
    – 7 Processes all doing final fit on ~500,000 molecule list (about 1 molecule/per sec!  These are gonna have to get killed…)
  • Finish O/N STORM run of 5E/AbdB data.
    – Still running at 9A, still focused.  will kill when Colenso is ready.
  • Check data downloads from Flybuzz
    – Server froze 25% of download complete (7.2GB).
    – Download resumed ~2PM
  • Check progress on Cajal.  Need to analyze 3 color hox data from Alistair4 (8/31/12) .
    – Cajal Hit a string of empty movies on en/Epc/K27 dataset 🙁
    – Movie 4 and on all blank /out of focus movies.  Deleted useless files
    – Cajal now analyzing both folders of 3 color hox data
  • Code improvements
    • Modify Cajal /Tuck run data to id all names lacking a _list.bin file and run those (deal with the whole error and abort thing).  Also modify code to skip conv images.  Updated in fxn_InsightM as well.
    • Jeff figured out script to tell matlab structure of bin file and import molecule list rapidly in one swoop.  Structure of bin file: ‘frame 0’ contains master molecule list (averaged molecule stats from all frames it was on).
  • Deleted empty movies 5-13 in 09/16 AbdA dataset.  Makes enough room to copy last nights run of E5 to Alistair6
  • Confocal imaging of 7/11/12 data. Slide 4 (actually sample 6: AbdA, AbdB, K27me3-647 no 750 stain).  Slide 6 (actually sample 5:
  • send slides to Mounia (fedex)
  • Discussions with Colenso and Yari: 488 switches off (not bleached, just dark state) in resin easily.  Suspect lack of 488 signal is switched off dyes.  This is why it looks like I can still see them in the STORM images and the confocal images in prolong Gold, but not the conventional images in STORM buffer.  Cy3B and 647 don’t switch off in resin much at all in comparison.  Try Alexa 514 or Alexa 532
  • Run im_nucseg on all unprocessed snail mRNA data.
  • Set up STORM analysis of 5E data on Monet (small run in STORM folder) and Cajal (O/N run in STORM2 folder).
  • Cut  4 coverslips of sample 4 from 7/11/12 (labeled as slide 6 in database).
  • Cut 3 coverslips of oligo-STORM sample (OS) from 7/11/12 collection.
This entry was posted in Summaries and tagged , , . Bookmark the permalink.